### v2

@@ 4.5 Ablation Studies
E-$L_{1}$ E-$L_{2}$ LCH Prompt 0.5 Avg. 0.5 Avg. 0.5 Avg. #params
@@
Table 4: Ablation study on the effect of audio-visual fusion for TAL and SED tasks (Tab. 4(a)) and different visual embeddings for all three tasks (Tab. 4(b)).
(a) “V” denotes visual-only, “A” denotes audio-only, and “A&V”denotes both audio and visual.
TAL Modality 0.5 0.75 0.95 Avg.
V 55.0 35.0 4.8 34.2
A&V 56.6 35.4 5.1 35.3
SED Modality 0.5 0.7 0.9 Avg.
A 51.4 38.1 13.9 49.5
A&V 61.0 45.9 19.8 57.7
(b) “FT” denotes using the visual encoder of ONE-PEACE fine-tuned on Kinetics-400.
TAL AVEL SED
Model FT 0.5 Avg. 0.5 Avg. 0.5 Avg.
ST 50.5 31.9 51.4 48.4 58.7 55.9
ST 56.6 35.3 53.2 49.6 61.0 57.7
AT 51.7 32.6 50.7 48.6 61.8 59.3
AT 57.1 36.1 54.1 50.7 64.0 59.8
4.6 Visualization and Discussion
Qualitative results. In Fig. 3, we visualize the predictions of our AT model on three tasks. Compared with the other state-of-the-art methods, our model generates localizations that better overlap with ground truth. For instance, our model outputs finer boundaries to distinguish repeated occurrences of the same events, e.g., the audio-visual event of “vacuum cleaner cleaning floors” and the sound event of “speech”. For SED, visual information helps the model detect the “frying” sound event more accurately compared to that using audio alone.
Figure 4: Examples of localizing the visual/sound events that are not present in ActivityNet 1.3/DESED datasets. The videos are from the test set of UnAV-100. “GT (AVEL)” denotes only AVEL annotations provided during training. “Pred” is the prediction results on each task. “score” is the confidence score of predictions.
Localizing instance categories across tasks. Benefiting from the language-aware classification head that utilizes the pre-trained text encoder [49], our AT model has a novel capability of localizing the categories from other task datasets when performing the current task. For example, we select some categories from the UnAV-100 dataset that are not present in ActivityNet 1.3 and DESED, and then tokenize them with the TAL and SED prompts and concat with the original class embeddings to conduct inference on all three tasks. In Fig. 4, we observe that our AT model successfully detects the visual events of “driving motorcycle”, “car passing by”, and “playing cornet”, etc., and the sound events of “church bell ringing” and “man speaking", etc., even though the model just learned the audio-visual events of these categories during training. It indicates that our model is capable of detecting all three types of instances (visual, sound and audio-visual) for the categories of all three datasets (total 310 classes).
Figure 5: Examples of open-vocabulary localization. The multi-task model trained on ActivityNet 1.3 for TAL and DESED for SED is utilized to detect unseen instance categories from UnAV-100. The videos are from the test set of UnAV-100.
Emergent open-vocabulary localization. We also explore our model’s ability of open-vocabulary localization. We use the multi-task model (row 4 in Tab. 1) trained on ActivityNet 1.3 and DESED and test it on the category set of UnAV-100. We visualize two examples in Fig. 5. There is no class “bull bellowing” and “playing drum kit”, even a similar one in ActivityNet 1.3 and DESED, but we surprisingly found that the model can accurately detect the corresponding visual and sound events with high confidence scores. This demonstrates that our model has strong potential in open-set capabilities. More ablation studies and qualitative examples can be found in the Supp. materials.
5 Conclusion
We propose a Unified Audio-Visual perception network (UniAV) for the joint learning of TAL, SED and AVEL tasks for the first time, realizing the localization of visual actions, sound events and audio-visual events in an untrimmed video by a single unified model. Specifically, we introduce a unified audio-visual encoding pipeline to minimize data discrepancies, while applying task-specific experts to capture distinct knowledge for each task. Moreover, a unified language-aware classifier allows the model to have high flexibility and generalizability during inference. Extensive experiments demonstrate that UniAV achieves superior and competitive performances on three challenging benchmarks, surpassing its single-task counterparts by a large margin with fewer parameters.
Limitations. As the first attempt, our model was trained on limited data and exhibits partial generalizability. In the future, we plan to utilize more available data and fully leverage existing large-scale multi-modal pre-trained models to further explore the model’s capabilities in open-world predictions.
Appendix 0.A More Implementation Details
We utilize the visual and audio encoders of ONE-PEACE [49] to tokenize visual and audio modalities for all three tasks, respectively. Specifically, each frame is first resized to ensure that its shortest side is 256 pixels, followed by cropping the center region to obtain 256 $\times$ 256 frames. Then, the RGB stacks are passed through the visual encoder, and the average pooling is applied on the temporal axis, producing a 1536-d feature for each stack of 16 frames. Also note that the visual encoder was fine-tuned on the Kinetics-400 [7] with inputs of 16 frames. For the audio modality, we directly input the waveform of 1s audio clips into the audio encoder to obtain a 1536-d feature for each clip. For our unified framework, in the audio-visual pyramid transformer, the number of attention heads is 4 in both uni-modal and cross-modal blocks. The temporal downsampling operation is realized by using a single depth-wise 1D convolution. Our models are trained on a Nvidia Tesla V100 GPU, and the code implementation relies on PyTorch framework and will be released upon publication.
Dataset selection for multi-task learning. Facing the fact that the existing datasets for untrimmed video localization tasks are limited, we chose ActivityNet 1.3 [6], UnAV-100 [11] and DESED [20] as they are currently the most mainstream datasets with relatively high quality and suitable for our multi-task model training. For TAL task, there are several other popular datasets, such as THUMOS14 [18] and EPIC-Kitchens 100 [9]. However, THUMOS14 contains very limited videos (200 videos for training) with only 20 uncommon sport classes, e.g., “CleanAndJerk”, “PoleVault” and “JavelinThrow”, etc. Besides, EPIC-Kitchens 100 consists of egocentric videos constrained to fine-grained kitchen scenarios. Thus, there exist extremely large domain gaps with the datasets for other tasks. By contrast, ActivityNet 1.3 has 200 rich common human activities including the top-level categories of housework, working, sports, eating and drinking, and animal caring, etc., which has domain overlaps with other task datasets, facilitating the model to learn mutually beneficial knowledge across tasks during multi-task training. Furthermore, only UnAV-100 dataset for AVEL task is based on untrimmed videos, and DESED is the only dataset whose audio and visual modalities are both available for SED task to the best of our knowledge.
Appendix 0.B More Results Analysis
Effect of parameter quantity. From the ablation study on the proposed task-specific experts in Sec. 4.5, we can see that adding experts on more transformer blocks leads to a significant increase in the number of parameters. In order to explore the effect of parameter quantity on our model, we reduce the dimension $D$ and $D^{\prime}$ of the transformer blocks. In Tab. 5, we can see that when the dimension drops to 352 with the parameter number halved, our unified all-task (AT) model still keeps superior performances with minor variations. And increasing parameters may not necessarily improve performances, i.e., increasing to 92M causes drops in SED results. Overall, it clearly proves that the performance improvement of our AT model is not due to an increase in parameters but the effectiveness of our proposed task-specific experts.
Table 5: Study on the effect of parameter quantity of our AT model. “Dimension $D/D^{\prime}$” denotes the dimension of the transformer blocks in the model.
Dimension $D/D^{\prime}$ TAL AVEL SED
0.5 Avg. 0.5 Avg. 0.5 Avg. # params
352 56.7 36.0 53.4 50.3 63.8 60.2 63M
424 57.3 36.1 54.6 50.8 62.4 59.7 92M
512 57.1 36.1 54.1 50.7 64.0 59.8 130M
@@ Appendix 0.B More Results Analysis
Table 6: Comparison with our models using I3D [7] and VGGish [16] features. “SOTA”: TriDet [42] for TAL, UnAV [11] for AVEL and SED.
@@
Table 7: Study on different text encoders for category embedding.
Text Encoder TAL AVEL SED
@@
RoBERTa [31] 45.5 28.8 43.9 41.5 61.3 57.9
CLIP [39] 55.7 34.7 53.5 49.8 62.0 58.1
ONE-PEACE [49] 57.1 36.1 54.1 50.7 64.0 59.8
Table 8: Study on different audio and visual encoders. We show the results on the public evaluation set of DESED for SED task.
@@
Experiments using I3D [7] and VGGish [16] features. We also evaluate our model using I3D [7] and VGGish [16] features that are commonly used on previous models. The results are shown in Tab. 6. Note that the I3D visual and VGGish audio embeddings are not aligned with the ONE-PEACE [49] text embeddings, which severely limits the effectiveness of our unified language-aware classifier (LCH). Alternatively, our AT model without LCH achieves performance boosts on all three tasks compared to our single-task (ST) models, indicating the great benefit of multi-task learning and the proposed task-specific experts. Besides, finetuning gains further performance boosts, setting new state-of-the-art results compared with existing methods with the same features. In conclusion, the lack of modality alignment (audio-visual-language) and limited generalization of traditional I3D and VGGish encoders constrain our model’s capability and flexibility. It clearly proves the great necessity of using a general model pre-trained by aligning vision, audio and language modalities (e.g., ONE-PEACE) as our audio/visual encoders.
Effect of different text encoders. In Tab. 7, we compare the performances of our multi-task (AT) model using different text encoders for category embedding in the language-aware classifier. RoBERTa [31] is a large natural language processing (NLP) model widely used in various language understanding tasks, and CLIP [39] is a visual-language pre-trained model learned from a vast data of image-text pairs. We can see that using RoBERTa, a purely NLP model, yields poor results across all three tasks, while due to the exceptional text encoding capability of the CLIP, significant performance boosts can be observed. Furthermore, when utilizing the text encoder of ONE-PEACE, our AT model achieves the best results on all three tasks. It indicates that our model can benefit from the modality-aligned representations by utilizing the visual, audio and text encoders from the same ONE-PEACE model.
@@
Effect of different audio and visual encoders. In Tab. 8, we also explore the performances of CLAP [51] and CLIP [39] for the encoding of audio and visual modalities, respectively. We show the results of the single-task model for the SED task. CLAP [51] is a popular language-audio pre-trained model for various downstream tasks, such as text-audio retrieval and audio classification. However, we find that using the audio embeddings extracted from CLAP leads to poor performance on the SED task. It may be attributed to its emphasis on global representations of long audio clips and insufficient fine-grained audio information modeling, making it unsuitable for localization tasks. Besides, we can see that the video features extracted from the CLIP [39] visual encoder have comparable performances with those from ONE-PEACE. When using both audio and visual modalities, the ONE-PEACE features can achieve the best results.
Appendix 0.C More Visualization Examples
Qualitative results. We visualize more prediction results of our all-task (AT) model on three tasks in Fig. 6, where we compare with the same methods and use the same setting as in Sec. 4.6. We can observe that our model obtains relatively better predictions than other state-of-the-art methods.
Localizing instance categories across tasks. More examples are presented in Fig. 7, where we use the same setting as in Sec. 4.6. Our AT model can successfully detect the visual and sound events of “tap dancing”, “woman speaking”, “female singing”, “lions roaring”, “man speaking” and “chainsawing trees”, even though the model just learned the audio-visual events of these categories during training. It further indicates that our AT model has a strong capability of detecting all three types of instances for the categories of all three datasets.
Emergent open-vocabulary localization. We also provide more examples in Fig. 8, where we use the same setting as in Sec. 4.6. There is no class of “playing tabla”, “airplane flyby”, “hammering nails”, and “helicopter”, even a similar one in ActivityNet 1.3 and DESED datasets, but we can see that the model can accurately detect the corresponding visual and sound events with relatively high confidence scores. It demonstrates our model’s potential capability of open-vocabulary localization.
Figure 6: More qualitative results on TAL, AVEL and SED tasks.
Figure 7: More examples of localizing instance categories across tasks.
Figure 8: More examples of open-vocabulary localization.

### v3

@@ IV-C Results and Analysis
$L_{1} E$-$L_{2}$ LCH Prompt 0.5 0.75 0.95 Avg. 0.5 0.6 0.7 0.8 0.9 Avg. 0.5 0.7 0.9 Avg. #params
1 55.4 34.9 6.3 34.8 53.0 47.4 41.0 33.5 21.5 49.8 62.7 50.5 26.2 58.6 62M
2 55.8 35.2 5.2 35.0 52.9 47.7 41.6 33.5 21.3 49.9 63.5 49.8 27.3 59.7 64M
3 56.9 36.1 5.7 35.9 53.7 48.5 42.5 33.9 20.5 50.5 61.4 49.8 26.4 58.5 80M
4 56.8 36.1 6.2 35.9 54.5 49.2 42.5 34.6 20.8 50.8 62.9 49.3 27.2 59.3 114M
5 56.2 35.7 5.1 35.2 54.6 49.4 42.4 34.6 22.6 51.2 57.4 43.5 19.5 54.8 133M
6 57.0 36.3 5.2 35.9 54.2 49.6 42.8 34.8 21.3 51.0 63.7 49.8 26.6 59.7 130M
7 57.1 36.4 5.5 36.1 54.1 48.6 42.1 34.3 20.5 50.7 64.0 49.5 27.0 59.8 130M
IV-D Comparison with Existing Methods
Table III presents the comparison results of our model with state-of-the-art works. We highlight that our UniAV (OursAT) is the only model that performs all three tasks by a single model parameters, while all other models in Table III are task-specific models that are separately trained for each task. For TAL task, we note that many previous TAL methods achieved superior results on ActivityNet 1.3 by combining with the external action classifier [76, 77]. By contrast, our model has good capabilities in both classification and regression and does not need to rely on any external classification models. For a fair comparison, we list the results of methods without using external classifiers in Table III. We simply concatenate audio and visual features as input for those TAL methods [9, 27, 8, 75, 28] when using both modalities. We can see that, with ONE-PEACE [21] features, our all-task (AT) model reaches an average mAP of $36.1\%$, outperforming other state-of-the-art task-specific models. We also trained UnAV [15] that is tailored for AVEL task to conduct TAL task, but found the result is much lower than highly specialized TAL models [27, 8]. It indicates that the task-specific models cannot be generalized effectively to other tasks, while our unified model can achieve superior or on-par performances on all three tasks with good generalizability. For AVEL task, our AT model gets very competitive results compared to UnAV [15] when using the same ONE-PEACE features. Besides, our single-task fine-tuned model (OursAT→ST) further improves the average mAP to $51.7\%$, setting a new state-of-the-art result on the AVEL task. For SED task, since the traditional SED methods only support super fine-grained sound spectrograms as input with the evaluation metric not suitable for our unified approach, we implemented Actionformer [8], TriDet [27] and UnAV [15] to conduct SED task. We can see that our AT model outperforms all other methods by a large margin. Overall, we emphasize that our proposed AT model (UniAV) holds superior or on-par performances on all three tasks using only a single unified model, and achieves the best All Tasks Average score compared with other methods, which significantly distinguishes our model from other task-specific models.
IV-E Ablation Studies
Effect of task-specific experts (TE) and unified language-aware classification head (LCH). First, we explore the effect of our TE on different layers of the pyramid transformer. As shown in Table IV, applying TE in the later $L_{2}$ cross-modal transformer blocks (row 4) has better results than that in the early $L_{1}$ uni-modal transformer blocks (row 3). It indicates that the later stages of cross-modal fusion can capture distinct knowledge that is more beneficial for each task. Besides, adding experts in all transformer blocks (row 7) can achieve the best performances on TAL and AVEL tasks. Second, we find that using separate task-specific classifiers (row 5) leads to quite unstable results. In contrast, notable improvements are observed on TAL and SED when using LCH with only categories as text tokens (row 6), especially boosting SED with a $4.5\%$ increase at average mAP. It could be attributed to the LCH based on the large language encoder [21], which enhances generalization and effectively avoids overfitting on the small dataset. Overall, combining both modules yields complementary effects, significantly improving all tasks (row 1 vs. row 7), with greater performance gains at lower thresholds.
TABLE V: Ablation study on the effect of parameter quantity. “TE” is short for task-specific experts. $D\&D^{\prime}$” denotes the dimension of the transformer blocks.
TE $D\&D^{\prime}$ TAL AVEL SED
0.5 Avg. 0.5 Avg. 0.5 Avg. # params
@@ IV-E Ablation Studies
TABLE VI: Ablation study on our models using I3D [60] and VGGish [74] features. “SOTA”: TriDet [27] for TAL, UnAV [15] for AVEL and SED using same features.
@@
TABLE VII: Ablation study on our models using ImageBind [78] and LanguageBind [79] as encoders. “TE” and “LCH” denote the task-specific experts and the language-aware classification head, respectively.
Encoder Method TAL AVEL SED
@@
ImageBind [78] ST 49.3 30.8 46.4 43.6 55.4 51.8
Base 48.4 30.2 45.9 43.7 56.0 52.2
w/ TE 49.4 31.3 48.8 46.0 56.5 54.0
w/ LCH 49.1 31.2 46.8 44.6 56.6 53.4
OursAT 50.6 31.8 48.7 45.8 57.1 54.1
LanguageBind [79] ST 50.6 31.4 48.2 44.8 57.0 54.0
Base 49.8 31.0 47.5 44.2 57.8 54.6
w/ TE 51.3 32.4 50.4 47.3 59.5 56.8
w/ LCH 50.9 32.1 49.0 46.1 59.6 56.8
OursAT 53.0 33.5 50.3 47.5 60.1 57.3
TABLE VIII: Ablation study on the effect of audio-visual fusion for TAL and SED tasks. “V” denotes visual-only, “A” denotes audio-only, and “A&V”denotes both audio and visual.
TAL Modality 0.5 0.75 0.95 Avg.
V 55.0 35.0 4.8 34.2
A&V 56.6 35.4 5.1 35.3
SED Modality 0.5 0.7 0.9 Avg.
A 51.4 38.1 13.9 49.5
A&V 61.0 45.9 19.8 57.7
TABLE IX: Ablation study on the importance of motion information during visual tokenization. “FT” denotes using the visual encoder of ONE-PEACE fine-tuned on Kinetics-400.
TAL AVEL SED
Model FT 0.5 Avg. 0.5 Avg. 0.5 Avg.
ST 50.5 31.9 51.4 48.4 58.7 55.9
ST 56.6 35.3 53.2 49.6 61.0 57.7
AT 51.7 32.6 50.7 48.6 61.8 59.3
AT 57.1 36.1 54.1 50.7 64.0 59.8
TABLE X: Ablation study on other large-scale audio and visual encoders. We show the results on the public evaluation set of DESED dataset for SED task.
@@
Effect of parameter quantity. From the study of the proposed task-specific experts (TE) in Table IV, we can see that adding experts on more transformer blocks leads to a significant increase in the number of parameters. In Table V, we explore the effect of parameter quantity on our model. We can observe that when removing the task-specific experts and increasing the dimensions $D$ and $D^{\prime}$, the model’s performance does not improve but declines across the three tasks. It clearly proves that the performance improvement of our model is not due to an increase in parameters but the effectiveness of our proposed task-specific experts.
@@
Effect of I3D and VGGish features. We also evaluate our model using I3D [60] and VGGish [74] features that are commonly used on previous models. The results are shown in Table VI. Note that the I3D visual and VGGish audio embeddings are not semantically aligned with the used ONE-PEACE [21] text embeddings, which severely limits the effectiveness of our unified language-aware classifier (LCH). Alternatively, our AT model without LCH achieves performance boosts on all three tasks compared to our single-task (ST) models, indicating the great benefit of the proposed task-specific experts. Besides, fine-tuning gains further performance boosts, setting new state-of-the-art results compared with existing methods with the same features. In conclusion, the lack of modality semantic alignment (audio-visual-language) and limited generalization of the traditional I3D and VGGish encoders constrain our model’s capability and flexibility. It clearly proves the great necessity of using a general model pre-trained by aligning vision, audio and language modalities (e.g., ONE-PEACE) as our audio/visual encoders.
Effect of other tri-modal aligned encoders. We also demonstrate that our model remains effective when using other existing tri-modal aligned encoders, including ImageBind [78] and LanguageBind [79]. We apply the same sampling strategy used for ONE-PEACE to extract features for a fair comparison. As shown in Table VII, our AT model consistently outperforms the Naive Multi-Task (Base) model across all three tasks when using both encoders. Besides, we can see that the proposed task-specific experts (TE) and the language-aware classification head (LCH) maintain their remarkable effectiveness. Therefore, our model does not rely on the specific encoder [21], rather, the semantically well-aligned embeddings are what support the effectiveness and flexibility of our model. We note that since the architecture and training strategy of ONE-PEACE [21] uniquely emphasize modality interaction and fine-grained understanding, along with the visual encoder fine-tuned on Kinetics-400 [60], it leads to superior encoding performance compared to the other tri-modal encoders.
Audio-visual fusion for TAL and SED. We also verify the effectiveness of audio-visual fusion for TAL and SED tasks in Table VIII. We use the single-task ST model (row 1 in Table I) for each task to conduct the experiments. For TAL, we can see the performance increase ($1.1\%$ at average mAP) as we insert audio signals to apply cross-modal interactions. For SED, the model obtains a substantial performance boost ($8.2\%$ at average mAP) when adding visual modalities, indicating the critical role of both modalities for the SED task.
Importance of motion information. In Table IX, we compare the performances of our models using different visual encoders. For both single-task (ST) and all-task (AT) models, the performances improve by a large margin when using the visual encoder [21] fine-tuned on Kinetics-400 [60]. In particular, for TAL, the improvements of $3.4\%$ and $3.5\%$ at average mAP can be observed for the ST and AT models, respectively. This emphasizes the importance of motion information for video localization tasks, especially for the TAL task.
@@
Comparison with other large-scale audio and visual encoders. In Table X, we also explore the performances of other large-scale audio and visual encoders. We show the results of the single-task model for the SED task. We find that using the audio embeddings extracted from CLAP [80] leads to poor performance. This may be attributed to its emphasis on global representations and insufficient fine-grained audio information modeling. Besides, CLIP [65], as a powerful visual-language pre-trained model, has comparable performances with other tri-modal aligned encoders. Moreover, LanguageBind [79] outperforms ImageBind [78] in audio feature representation but is slightly weaker in visual features. When using both audio and visual modalities, the ONE-PEACE features can achieve the best results.
Comparison with other text encoders. In Fig. 5, we compare the performances of our all-task (AT) model using different text encoders for category embedding in the unified language-aware classifier. All models apply the audio and visual encoders of ONE-PEACE [21] for video representations by default. We can see that RoBERTa [81], a large natural language processing (NLP) model, yields poor results across all three tasks, while due to the exceptional text encoding capability of CLIP, significant performance boosts can be observed. Besides, ImageBind [78]’s text encoder performs better than that of LanguageBind [79] on TAL and SED tasks. Furthermore, when utilizing the text encoder of ONE-PEACE, our AT model achieves the best results on all three tasks.
Figure 5: Localization results of our AT model on all three tasks with different text encoders for category embedding, measured by mAP@tIoU=0.5.
Figure 6: Examples of localizing the visual/sound events that are not present in ActivityNet 1.3/DESED. The videos are from the test set of UnAV-100. “GT (AVEL)” denotes only AVEL annotations provided during training. “Pred” is the prediction results on each task. “score” is the confidence score of predictions.
Figure 7: t-SNE visualization of the pyramid feature $Z^{1}$ learned by (a) UniAV with task-specific experts and (b) the multi-task variant without task-specific experts. Different colors denote different tasks, and each point denotes the feature of a temporal segment from the given video.
Figure 8: t-SNE visualization of (a) the class embeddings with task prompts encoded by ONE-PEACE text encoder, and (b) the last FC layer weights from the separate task-specific classification heads. Different colors denote different tasks, and each point denotes an embedding of an instance category.
IV-F Visualization and Discussion
Qualitative results. In Fig. 4, we visualize the predictions of our AT model on three tasks. Compared with the other state-of-the-art methods, our model generates localizations that better overlap with the ground truth. For instance, our model outputs finer boundaries to distinguish repeated occurrences of the same events, e.g., the audio-visual event of “vacuum cleaner cleaning floors” and the sound event of “speech”. For SED task, visual information helps the model to detect the sound event of “frying” more accurately compared to that using audio modality alone.
t-SNE visualization for TE and LCH. We first randomly sample a video for inference and visualize the output temporal feature sequences from the audio-visual pyramid transformer (level 1). As shown in Fig. 7 (a), using task-specific experts (TE), the features for different tasks are clearly distinguishable, while without TE (Fig. 7 (b)), the features almost overlap. This reveals that TE enables the model to effectively learn distinct knowledge for each task. Next, we visualize the class embeddings encoded by ONE-PEACE for unified classification (Eq.6). As in Fig. 8 (a), the embeddings reveal rich correlations between event categories across tasks, e.g., “a visual event of {playing drums}” and “an audio-visual event of {playing drum kit}”. In contrast, the last FC layer weights of separate task-specific classification heads show no inter-task correlations (Fig. 8 (b)). It suggests that our LCH breaks the constraint of fixed, independent class sets of different tasks, effectively fostering semantic associations between events across tasks.
Localizing instance categories across tasks. Benefiting from the language-aware classification head that utilizes the pre-trained text encoder [21], our AT model has a novel capability of localizing the categories from other task datasets when performing the current task. For example, we select some categories from the UnAV-100 dataset that are not present in ActivityNet 1.3 and DESED, and then tokenize them with the TAL and SED prompts and concatenate with the original class embeddings to conduct inference on all three tasks. In Fig. 6, we observe that our AT model successfully detects the visual events of “driving motorcycle”, “car passing by”, and “playing cornet”, etc, and the sound events of “church bell ringing” and “man speaking”, etc, even though the model just learned the audio-visual events of these categories during training. It indicates that our model has the remarkable flexibility to localize rich instance categories across tasks by simply changing task prompts, greatly alleviating the data scarcity for each task, especially for SED. UniAV is capable of detecting all three types of instances (visual, sound and audio-visual) for the categories of all three datasets (total 310 classes).
Figure 9: Examples of open-vocabulary localization on UnAV-100. The multi-task model trained on ActivityNet 1.3 for TAL and DESED for SED is utilized to detect unseen instance categories from UnAV-100. The videos are from the test set of UnAV-100.
Figure 10: Examples of open-vocabulary localization on LLP [40] and LFAV [44] using our UniAV trained on ActivityNet 1.3 (TAL), DESED (SED) and UnAV (AVEL). The videos are from the LLP (left) and LFAV (right) test sets, respectively. “GT” is the LLP/LFAV annotations of visual (V) and audio (A) events.
Emergent open-vocabulary localization. Apart from significant gains for closed-set performances, our model also showcases an impressive ability of open-set localization for novel categories, owing to the generalizability of the used tri-modal model [21] and the unified language-aware classification head. First, as shown in Fig. 9, there is no class “bull bellowing” and “playing drum kit”, even a similar one in ActivityNet 1.3 and DESED category sets, but we surprisingly found that the multi-task model (row 4 in Table I) can accurately detect the corresponding visual and sound events with high confidence scores. Additionally, we also conduct open-set localization on LLP [40] and LFAV [44] datasets. In Fig. 10, although our training data does not contain the categories “cello”, “Violin_fiddle”, and “fixed-wing_aircraft” (only the similar ones like “playing cello/violin” and “airplane flyby”). Despite this, our model successfully detects the corresponding visual/sound/audio-visual events with high confidence. It demonstrates our model’s strong potential in open-set capabilities, greatly distinguishing it from other approaches.
V Conclusion
This paper proposes a Unified Audio-Visual perception network, UniAV, for the joint learning of TAL, SED and AVEL tasks for the first time, realizing the localization of visual actions, sound events and audio-visual events in an untrimmed video by a single unified model. Specifically, we introduce a unified audio-visual encoding pipeline to minimize data discrepancies across tasks, and meanwhile applying task-specific experts to capture distinct knowledge for each task. Moreover, a novel unified language-aware classifier allows the model to have high flexibility and generalizability during inference. Extensive experiments demonstrate that our UniAV achieves superior and competitive performances on three challenging benchmarks, surpassing its single-task counterparts and the naive multi-task baseline using only a single model parameters. In the future, we plan to utilize more available data and fully leverage existing large-scale multi-modal pre-trained models to further explore the model’s capabilities in open-world predictions.